OpenStudy (anonymous):

i have another question quys: Graph the following function using transformations. Be sure to graph all of the stages on one graph. State the domain and range. For example, if you were asked to graph y= x2 + 1 using transformations, you would show the graph of y = x2 and the graph shifted up 1 unit. Please do not show only the final graph. the function is: y=x^2+14x

OpenStudy (anonymous):

so you want that graph shifter up 1 unit?

OpenStudy (anonymous):

idk thats y im asking yall lol

OpenStudy (anonymous):

well start with f(x)=x^2 then transform like this f(x)+14x = x^2+14x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

domain is infinity

OpenStudy (anonymous):

range is [5,infinity)